OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1

OpenStudy (sidsiddhartha):

\[a_1=-3 ~~~~~and~~~~~~~~a_n=2*a_{n-1} \\right?\]

OpenStudy (sidsiddhartha):

first try putting n=2 in\[a_n=2*a_{n-1}\] so u'll have \[a_2=2*a_{2-1}=2*a_1=2*(-3)=-6\] now put n=3 then \[a_3=2*a_{3-1}=1*a_2=2*(-6)=-12\] after that try n=4 , then n=5 and u'll get the answer